OpenStudy (anonymous):

Insert 19 A.M. between -1/4 and -9 3/4.

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

the ans. is -1/4 , -3/4,..........,-9 3/4

OpenStudy (anonymous):

if you want to insert n arithmetic means between a and b, you can add this number to a untill reach b. the number : d = (b-a)/ n+1 find d

OpenStudy (anonymous):

i have tried but not getting the ans.

OpenStudy (anonymous):

ok..i ll try

OpenStudy (anonymous):

and

OpenStudy (anonymous):

in my book it is printed 1/4 and not -1/4. but from ans. i've posted it -1/4

OpenStudy (anonymous):

ya if you start from 1/4 as a and -9 3/4 as b you get the series as -1/4,-3/4,.........

OpenStudy (anonymous):

d= -17/20 ?

OpenStudy (anonymous):

if a=1/4 and b=-9 3/4 ie b=-39/4 d=(-39/4-1/4)/(20) d=-1/2

OpenStudy (anonymous):

b=-39/4 or -33/4?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b=-39/4

OpenStudy (anonymous):

sorry earlier was by mistake...

OpenStudy (anonymous):

what is b?

OpenStudy (anonymous):

-39/4??

OpenStudy (anonymous):

b=-9 3/4 b=((-9)(4)+3)/4 b=(-36+3)/4 b=-39/4

OpenStudy (anonymous):

-36+3 is -33 ??

OpenStudy (anonymous):

i think it is -(9+3/4)

OpenStudy (anonymous):

see just forget negative sign for a while./...... 9 3/4=((9*4)+3)/4=(36+3)/4=39/4 now put negative sign.....

OpenStudy (anonymous):

and if you divide 33/4 you dont get 9 3/4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ya it means same -(9+3/4)=-(39)/4

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

anything else......??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait